OpenStudy (anonymous):

two to the 3rd power *2 to the negative5 power*2 to the fifth power

OpenStudy (anonymous):

\[2^{3} * 2^{-5} * 2^{5} = 2^{3-5+5} = 2^{3} = 8\]

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

2^3*2^-5*2^5........^=to the power of 2^3+-5+5=2^3=8

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

nice amanda

OpenStudy (anonymous):

thanks..is it good?

OpenStudy (anonymous):

i think so. im in advanced so probably

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

>_<tehetehe

OpenStudy (anonymous):

next question?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

sure u pick this tme

OpenStudy (anonymous):

ok just look for my chicken picture

OpenStudy (anonymous):

k